Find the slope and y-intercept of the line represented by the following equation.
2x - 3y = 6

Answer by ptaylor(2198)


The standard form for slope and y-intercept is:
y=mx+b where m is the slope and b is the y intercept
Now lets put your equation into standard form:
2x - 3y = 6 first, subtract 2x from both sides
2x-2x-3y=6-2x or
-3y=-2x+6 next, divide both sides by -3
y=(2/3)x-2
slope (m)=2/3
y-intercept (b)=-2

Hope this helps----ptaylor